OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8 A. y== -3/2x+2 B. y=-3/2x+4 C. y= 3/2x+4 D. y=3/2x+2 D. 07_07_LQ_1a.gif

OpenStudy (anonymous):

I think is a

OpenStudy (shinalcantara):

Slope-intercept form y = mx + b where m is the slope and b is the y-intercept

OpenStudy (shinalcantara):

Given the equation 6x + 4y = 8 you'll have: ------------------ 6x + 4y = 8 Transpose 6x to the right side of the equation 4y = -6x + 8 Divide the whole equation with 4 4y/4 = -6x/4 + 8/4 y = -6x/2 + 2 y = -3/2 x + 2